# Entropy-Gated Branching for Efficient Test-Time Reasoning

### v2

@@ Abstract
Test-time compute methods like beam search can significantly improve the reasoning capabilities and problem-solving accuracy of large language models. However, these approaches require substantially increased computational resources, with most computation wasted on exploring low-diversity branches where the model already exhibits high confidence. We observe that a small subset of uncertain reasoning steps has a disproportionately large impact on final prediction accuracy, and branching at these points tends to yield higher-quality and more diverse candidate reasoning steps. Therefore, we introduce Entropy‑Gated Branching: a novel inference technique that dynamically allocates computational resources by selectively expanding prediction sequences only at points of high uncertainty. Our method leverages entropy as a gating mechanism to identify when branching is most beneficial, coupled with an external feedback model to rank and prune candidate branches. Empirical results on mathematical and financial reasoning benchmarks show that this strategy improves accuracy by 22.6% over standard inference while operating 37% faster than conventional beam search with similar or higher performance. Our results show that dynamic resource allocation during inference can substantially improve both efficiency and effectiveness, offering a more scalable pathway to enhanced LLM reasoning capabilities.
@@ 1 Introduction
Language models have demonstrated remarkable capabilities across a diverse range of tasks (Bang et al. 2023; Li et al. 2023; Mahfouz et al. 2024), yet persist in making fundamental errors due to hallucinations or flawed reasoning (Huang et al. 2024; Tong et al. 2024; Farquhar et al. 2024; Zhang et al. 2025; Galitsky 2025; Penny-Dimri et al. 2025). These mistakes are especially detrimental in high-stakes domains such as finance and healthcare. To address this issue, it is crucial to develop models that ensure outputs are both verifiable and interpretable, reducing the risk of critical mistakes.
@@
Consequently, understanding the underlying causes of these errors and devising practical mitigation strategies is vital for building more reliable language models. Many studies (Xie et al. 2025; Yu et al. 2025; Wei et al. 2025) have demonstrated that fine-tuning language models using rewards or high-quality data can significantly enhance their performance. However, these approaches require substantial time and computational resources to retrain the model. Recently, the research community has shifted focus to test-time compute (TTC) as a complementary approach to improving model performance at low cost (Bi et al. 2024; Liang et al. 2024b; Beeching, Tunstall, and Rush 2024; Snell et al. 2025). TTC methods offer the compelling advantage of enhancing model capabilities without modifying trained weights, enabling performance scaling at inference time. Most TTC methods operationalize extra computation through extensive search algorithms, expanding multiple reasoning branches at every step and selecting the most promising continuation.
@@
However, indiscriminately exploring all candidate branches at every step is very time-consuming. Test-time search algorithms could take minutes to explore all potential intermediate reasoning chains before generating the final answer. This is not user-friendly and poses a significant barrier to real-world deployment. To understand when that extra test-time computation is worthwhile, we analyzed the entropy of the next-token logit distributions on mathematical reasoning benchmarks (Wang et al. 2025). As shown in Figure 1, models are considerably more error-prone during entropy spikes. Since a spike implies a broad posterior over tokens, allocating extra expansion precisely at these high-entropy steps yields maximal diversity and a better chance to rescue an otherwise faulty trajectory; conversely, when entropy is low, the model exhibits high confidence, making further expansion less fruitful since it would likely produce similar continuations. Empirically, we found that 39 out of 50 occurrences that coincide with a high-entropy spike lead to a flawed step of reasoning. These observations suggest that high entropy moments can indicate effective points to branch the model’s reasoning, rather than wasting computation by branching at every step. This raises the question: Is it possible to improve TTC efficiency by selectively expanding beams while still maintaining the performance benefits of beam search?
@@
We propose Entropy-Gated Branching (EGB), a novel TTC inference method that uses the entropy of the model’s predicted logit distribution to identify optimal branching points, instead of expanding at every step. We adopt entropy as a measure of uncertainty to concentrate search budget on high-impact uncertainty moments, while confident beams continue with standard sampling, avoiding unnecessary computational overhead. EGB demonstrates the ability to match or outperform test-time beam search, while achieving substantial reductions in inference time and computational costs.
@@ 2 Related Work
Handling uncertainty in language models is crucial for ensuring reliable decision-making and response generation. Prior work (Manakul, Liusie, and Gales 2023; Abbasi-Yadkori et al. 2024; Wagner et al. 2024) has explored sampling multiple responses to detect inconsistencies as a measure of uncertainty. However, despite their effectiveness, those methods cannot assess the reliability of a single response, require multiple iterations instead, making them inefficient for real-world applications. An alternative approach leverages token-probability methods (Fadeeva et al. 2024; Duan et al. 2024), which assess uncertainty at the token level without requiring multiple resampling steps.
@@
Recent research (Snell et al. 2025; Liu et al. 2025; Ji et al. 2025) has demonstrated that instead of allocating substantial GPU resources for training new models, enhancing computational resources during inference can yield significant performance gains. Currently, most test-time computing methods rely on generating multiple solutions through repeated sampling and then selecting the best answer. Many search algorithms like Monte-Carlo Tree Search (Zhou et al. 2023; Zhang et al. 2023; Liu et al. 2024a; Koh et al. 2024), guided beam search (Xie et al. 2024) or hybrid search (Snell et al. 2025) show that using search at inference time can lead to performance gains in various tasks. Another line of work (Muennighoff et al. 2025) appends a “wait” token as a method to control the budget and enhance a model’s reasoning process during inference.
LLM Feedback
In language models, feedback is often introduced to improve performance (Pan et al. 2024; Li 2025; Estévez-Ayres et al. 2025). Recent studies have explored various methods to align language models with human preferences (Liang et al. 2024a; Scheurer et al. 2023), modifying prompts (Zhou et al. 2022; Liu et al. 2024b; Ma et al. 2024) or refining errors (Jimichi, Funakoshi, and Okumura 2023; Kirstein, Ruas, and Gipp 2024). Integrating more sophisticated feedback loops will be essential for developing language models that are more reliable to diverse applications.
Figure 2: Illustration of the Entropy-Gated Branching: Left: The normal decoding where the model flows naturally. Middle: Traditional beam search samples KWKW candidates and depend on PRM score keep the top $K$ beams. Right: Entropy-gated branching expands uncertain beams at high entropy moments and generates confident beams normally.
@@ 3 Methodology
The principal limitation of this approach is its rigidly uniform search breadth. Exploring a fixed number of branches at every step often wastes significant compute when the model is confident and unable to produce diverse, high-quality alternate candidate sequences.
@@ 3.1 Entropy-Gated Branching
To improve computational efficiency, we propose Entropy-Gated Branching (EGB), a selective expansion strategy that allocates resources only where uncertainty warrants further exploration. Rather than expanding every beam uniformly, EGB uses the entropy of the model’s logit distribution as a dynamic gating mechanism. The entropy of a model’s logit distribution at timestep $t$ is defined as:
@@
where $p_{i,t}$ denotes the predicted probability of token $i$ at step $t$, and $V$ is the vocabulary size. Higher entropy corresponds to greater uncertainty and indicates a more uniform probability distribution over tokens.
We also calculate varentropy as shown in Figure 1, measuring the variance in entropy values over multiple potential outputs, defined as:
$V_{t}=\sum_{i=1}^{V}p_{i,t}\left(\log_{2}p_{i,t}+H_{t}\right)^{2}$ (2)
By capturing fluctuations in uncertainty, varentropy highlights situations where the model exhibits inconsistent confidence levels across tokens and helps us get a better understanding of the high-entropy moments.
This leads to two distinct cases for handling beams at each step, based on a pre-defined entropy threshold, $\tau$.
@@
When entropy exceeds the threshold, it indicates a critical decision point where the model is uncertain and multiple reasoning paths may be viable. These are the moments where exploration is most valuable. The uncertain beam is expanded into multiple diverse candidate continuations. For a beam $i$, this creates a set of $W$ candidate branches. This concentrates search budget precisely where it is needed most, increasing the likelihood of discovering a correct reasoning path that a normal decoding approach may miss.
@@ 3.2 Formalizing Selective Expansion
For uncertain beams requiring expansion, we employ multinomial sampling with temperature scaling to ensure diversity among the $W$ generated candidates. Specifically, for each uncertain beam $i\in\mathcal{U}t$, we sample from the probability distribution:
$p^{\prime}{j,t}=\frac{\exp(l_{j,t}/\tau_{temp})}{\sum_{k=1}^{V}\exp(l_{k,t}/\tau_{temp})}$ (4)
where $l_{j,t}$ represents the raw logit for token $j$ at timestep $t$, and $\tau_{temp}$ is a temperature parameter that controls the diversity of sampling. A lower temperature encourages more focused sampling around high-probability tokens, while higher temperatures promote greater exploration of the token space.
The total number of candidates generated at step $t$ is therefore at most $|\mathcal{P}_{t}|=K+(W-1)|\mathcal{U}_{t}|$. Since $|\mathcal{U}_{t}|\leq K$, this is always less than or equal to the $KW$ candidates produced by standard beam search. The computational complexity of selective expansion scales as $O(|\mathcal{U}_{t}|\cdot W\cdot V)$ for candidate generation, compared to $O(K\cdot W\cdot V)$ for uniform beam search. In practice, since $|\mathcal{U}_{t}|$ is typically much smaller than $K$ due to the selective nature of branching, EGB achieves substantial computational savings while maintaining search effectiveness.
@@ 3.3 Beam Evaluation
Before proceeding to quality assessment, EGB implements a duplicate detection mechanism to avoid wasting computational resources on scoring identical continuations. Since both uncertain beams and certain beams contribute to the candidate pool $\mathcal{P}_{t}$, identical or near-identical sequences can emerge from different sources, particularly when the model exhibits strong preferences for specific token sequences. Therefore, we employ string matching on the token-level and filter out the same branches first and gather the rest diverse branches for further evaluation. After collecting candidate continuations, assessing their quality is critical, as it directly influences the overall performance of the base model. A strong feedback model ensures that the generation follows a coherent and logical trajectory while minimizing errors. We employ a Process Reward Model that evaluates the quality of reasoning steps rather than just final outcomes. The PRM assigns scores to partial solutions based on their logical consistency, mathematical accuracy, and adherence to sound reasoning principles. For a candidate continuation $c_{t,j}^{(i)}$, the PRM score is computed as:
@@ 4 Experiments
CFA exams are renowned for their rigorous assessment of financial, quantitative, and analytical reasoning, making them an excellent proxy for evaluating model performance in high-stakes, real-world financial decision-making scenarios. Mistakes in the financial questions can lead to significant losses, reinforcing the need for precise and reliable reasoning. As official CFA exam questions are not publicly accessible, this study utilizes CFA mock exams purchased from AnalystPrep (AnalystPrep 2024), covering two levels of the CFA program. The dataset comprises multiple-choice and essay questions, each supplemented with corresponding answers, explanations, grading criteria, and metadata indicating the CFA topic associated with each question.
In addition to CFA data, we evaluate our methods on competition and general mathematical reasoning benchmarks of varying difficulty. First, we include an AIME evaluation set drawn from the AI-MO dataset (Art of Problem Solving Community 2025). This corpus contains all 90 problems from the 2022–2024 American Invitational Mathematics Examination (AIME I & II each year) (Art of Problem Solving Community 2025). Each AIME test is widely viewed as hard in difficulty between grade-school word problems and full Olympiad-style proofs (Mathematical Association of America 2025; Wikipedia contributors 2025).
We also assess performance on the MATH benchmark (Hendrycks et al. 2021), which consists of competition-level problems (AMC, AIME, and other sources) spanning a wide range of high-school mathematics topics and difficulty levels; for all experiments, we use 500 held-out test questions. Finally, we report results on GSM8K (Cobbe et al. 2021), a widely used grade-school math word-problem benchmark designed to probe multi-step arithmetic reasoning in language models, utilizing 1,319 test questions in our experiments.
@@
We implement our method on two families of models, Qwen3 (Yang et al. 2025) and Llama3 (Touvron et al. 2023), at 3 sizes each. For Qwen3 we use Qwen3-1.7B, Qwen3-4B, and Qwen3-8B, all in ”non-thinking” mode (Yang et al. 2025). For Llama3 we use Llama-3.2-1B-instruct, Llama-3.2-3B-instruct, and Llama-3.1-8B-instruct. We adopt the recommended sampling configuration for all experiments, with a temperature of 0.7, top-p of 0.8, top-k of 20, and min-p of 0 for Qwen3 with non-thinking mode and a temperature of 0.6, top-p of 0.9 for Llama models
@@
To balance performance with computational overhead, we set the entropy thresholds $\tau=2$ for AIME, CFA Level I and Level II math datasets, and $\tau=1.5$ for the GSM8K and MATH500 datasets since models exhibit less uncertainty on general math questions. For results reported in Table 1, we use a beam size $K$ of 3, a beam width $W$ of 5 for all datasets. Detailed analyses of the threshold values and branch numbers can be found in Section 5.3.
@@
All experiments were performed on A100 or H100 GPUs with 80GB VRAM, on Ubuntu 22, CUDA 12.4, Python 3.12. Each experiment was run on one GPU, utilizing 20-70 GB of GPU memory, depending on the model size and sequence length.
@@
Beam Search 54.83% 44.32% 11.11% 75.20% 84.00%
EGB 55.78% 43.18% 11.11% 75.00% 86.35%
@@
Beam Search 69.11% 56.57% 21.11% 84.00% 89.16%
EGB 70.89% 55.81% 18.60% 84.40% 90.67%
@@
Beam Search 76.12% 58.86% 18.89% 84.60% 88.25%
EGB 77.22% 58.52% 22.22% 82.20% 92.19%
@@
Beam Search 39.61% 42.58% 2.22% 39.40% 61.26%
EGB 34.09% 43.91% 2.22% 36.40% 54.89%
@@
Beam Search 49.06% 44.66% 6.67% 48.20% 82.18%
EGB 50.67% 45.45% 8.89% 51.40% 83.09%
@@
Beam Search 54.06% 47.73% 6.67% 47.60% 88.74%
EGB 57.50% 49.43% 8.89% 52.40% 88.10%
Table 1: Performance of EGB compared with beam search and standard decoding on finance and math benchmarks.
LLM-based Evaluation
Final answer correctness was determined using a powerful external language model, Gemini 2.5 Flash, as an impartial judge (Gemini Team et al. 2024). This approach circumvents the limitations of rigid evaluation methods like exact string matching or regex-based parsers, which are often brittle and can penalize semantically correct answers due to minor stylistic or formatting variations. For each problem, the LLM judge was provided with the original question, the model’s complete generated response, and the ground-truth solution, and was then prompted to assess whether the final answer was correct. This method ensures a robust and fair assessment by focusing on the semantic correctness of the reasoning outcome rather than superficial syntactic differences, providing a more reliable measure of the model’s true problem-solving capabilities.

### v3

@@ Abstract
Test-time compute methods can significantly improve the reasoning capabilities and problem-solving accuracy of large language models (LLMs). However, these approaches require substantially more computational resources, with most compute wasted on exploring low-diversity branches where the model already exhibits high confidence. We observe that a small subset of uncertain reasoning steps has a disproportionately large impact on final prediction accuracy, and branching at these critical junctures tends to yield more diverse and higher-quality candidate reasoning steps. We propose Entropy-Gated Branching (EGB), which branches only at high-uncertainty steps and prunes expansions with a lightweight verifier. On mathematical and financial reasoning benchmarks, EGB improves accuracy by 22.6% over standard inference while operating 31%-75% faster across math benchmarks than test-time beam search with higher performance. Our results show that dynamic resource allocation during inference can substantially improve both efficiency and effectiveness, offering a more scalable pathway to enhanced LLM reasoning capabilities.
Entropy-Gated Branching for Efficient Test-Time Reasoning
Xianzhi Li1, Ethan Callanan1, Abdellah Ghassel1, Xiaodan Zhu1, 1Queen’s University {li.xianzhi, e.callanan, abdellah.ghassel, xiaodan.zhu}@queensu.ca
@@ 1 Introduction
Language models have demonstrated remarkable capabilities across a diverse range of tasks Bang et al. (2023); Li et al. (2023); Mahfouz et al. (2024), yet persist in making fundamental errors due to hallucinations or flawed reasoning Huang et al. (2024); Tong et al. (2024); Farquhar et al. (2024); Zhang et al. (2025); Galitsky (2025); Penny-Dimri et al. (2025). These mistakes are especially detrimental in high-stakes domains such as finance. To address this issue, it is crucial to develop models that ensure outputs are both verifiable and interpretable, reducing the risk of critical mistakes.
@@
Consequently, understanding the underlying causes of these errors and devising practical mitigation strategies is vital for building more reliable language models. Many studies Xie et al. (2025); Yu et al. (2025); Wei et al. (2025) have demonstrated that fine-tuning language models using rewards or high-quality data can significantly enhance their performance. However, these approaches require substantial time and computational resources to retrain the model. Recently, the research community has shifted focus to test-time compute (TTC) as a complementary approach to improving model performance without modifying trained weights Bi et al. (2024); Liang et al. (2024b); Beeching et al. (2024); Snell et al. (2025). TTC methods offer the compelling advantage of enhancing model capabilities without modifying trained weights, enabling performance scaling at inference time. Many TTC methods operationalize extra computation through extensive resampling Wang et al. (2022), or searching multiple reasoning branches at every step and selecting the most promising continuation Xie et al. (2023); Snell et al. (2025).
@@
However, test-time search algorithms could take minutes to explore all potential reasoning chains before generating the final answer. This is not user-friendly and poses a significant barrier to real-world deployment. To understand when that extra test-time computation is worthwhile, we analyzed next-token logit entropy on mathematical reasoning benchmarks Wang et al. (2025). As shown in Figure 1, models are considerably more error-prone during entropy spikes. Since a spike implies a broad posterior over tokens, allocating extra expansion precisely at these high-entropy steps yields maximal diversity and a better chance to rescue an otherwise faulty trajectory; conversely, when entropy is low, the model exhibits high confidence, making further expansion less fruitful since it would likely produce similar continuations. Empirically, we found that 39 out of 50 occurrences that coincide with a high-entropy spike lead to a flawed step of reasoning. These observations suggest that high entropy moments can indicate effective points to branch the model’s reasoning, rather than wasting computation by branching at every step. This raises the question: Is it possible to improve TTC efficiency by selectively expanding beams while still maintaining the performance benefits of beam search?
@@
We propose Entropy-Gated Branching (EGB), a novel TTC inference method that uses the entropy of the model’s predicted logit distribution to identify optimal branching points, instead of expanding at every step. By adopting entropy as an uncertainty signal and combining it with external feedback models to jointly steer the search, EGB concentrates the budget on high-impact uncertainty moments, while confident beams continue with standard sampling, avoiding unnecessary computational overhead. EGB demonstrates the ability to outperform other test-time search algorithms while achieving substantial reductions in inference time and computational costs.
@@ 2 Related Work
Handling uncertainty in language models is crucial for ensuring reliable decision-making and response generation. Prior work has explored sampling multiple responses to detect inconsistencies as a measure of uncertainty Manakul et al. (2023); Abbasi-Yadkori et al. (2024); Wagner et al. (2024). However, despite their effectiveness, these methods cannot assess the reliability of a single response and require multiple iterations, making them inefficient for real-world applications. An alternative approach leverages token-probability methods Fadeeva et al. (2024); Duan et al. (2024), which assess uncertainty at the token level without requiring multiple resampling steps. Our work builds on token-level uncertainty methods by using entropy as a real-time signal to guide branching decisions during generation, rather than as a post-hoc uncertainty estimate. This allows EGB to dynamically allocate computation based on the model’s confidence at each step.
Diversity and Efficiency in Decoding
A fundamental challenge in generation is balancing diversity with computational efficiency. Standard beam search maintains multiple hypotheses but explores uniformly at each step, often generating redundant or low-quality alternatives Holtzman et al. (2019); Kulikov et al. (2018); Wang et al. (2022); Snell et al. (2025). Diverse beam search methods Vijayakumar et al. (2018) address this by explicitly encouraging dissimilarity among beams. Self-evaluation-guided stochastic beam search (SEGBS) Xie et al. (2023) introduced self-evaluation guidance via stochastic beam search, but still expands at every timestep regardless of model confidence. In contrast, EGB fundamentally differs from prior work by making branching itself uncertainty-aware rather than treating all generation steps equally.
@@
Recent research has demonstrated that enhancing computational resources during inference can yield significant performance gains without retraining models Snell et al. (2025); Liu et al. (2025); Ji et al. (2025). Most test-time computing methods rely on generating multiple solutions through repeated sampling and selecting the best answer. Search algorithms like Monte-Carlo Tree Search Zhou et al. (2023); Zhang et al. (2023); Liu et al. (2024a); Koh et al. (2024), guided beam search Xie et al. (2024), and hybrid search Snell et al. (2025) show that inference-time search can improve performance across various tasks. Another approach Muennighoff et al. (2025) appends "wait" tokens to control computational budget during reasoning. However, these methods typically apply uniform computational effort across all generation steps. EGB addresses this inefficiency by identifying when and where additional computation is most beneficial through entropy-based gating, achieving the performance benefits of test-time search with substantially reduced computational overhead.
Process Reward Models and Feedback
In language models, external feedback mechanisms are essential for guiding generation toward higher-quality outputs Pan et al. (2024); Li (2025); Estévez-Ayres et al. (2025). Recent work has explored aligning models with human preferences Liang et al. (2024a); Scheurer et al. (2023), prompt optimization Zhou et al. (2022); Liu et al. (2024b); Ma et al. (2024), and error refinement Jimichi et al. (2023); Kirstein et al. (2024). Process reward models (PRMs) have emerged as particularly effective for step-wise reasoning tasks, evaluating intermediate steps rather than only final answers Lightman et al. (2023); Uesato et al. (2022). EGB integrates PRMs as a critical component for ranking candidate branches generated at uncertainty spikes, ensuring that expanded beams are selected based on reasoning quality rather than arbitrary heuristics.
Figure 2: Illustration of Entropy-Gated Branching; Left: normal decoding where the model flows naturally.
Middle: traditional beam search samples $KW$ candidates and uses PRM scores keep the top $K$ beams.
Right: EGB expands uncertain beams at high entropy moments and generates confident beams normally.
@@ 3 Methodology
Entropy-Gated Branching (EGB) replaces this uniform policy with selective expansion. Instead of branching everywhere, EGB monitors uncertainty during decoding and only allocates extra branches at positions where the model is uncertain. Section 3.1 details the gating and rollback mechanism; Section 3.2 formalizes the resulting candidate pool and budget.
@@ 3.1 Entropy-Gated Branching
While decoding the next reasoning step for a beam, we compute the entropy of a model’s logit distribution at timestep $t$ as:
@@
where $p_{i,t}$ denotes the predicted probability of token $i$ at step $t$, and $V$ is the vocabulary size. High entropy indicates a flatter distribution and thus higher uncertainty.
The decision to branch is controlled by a crucial hyperparameter: the entropy threshold, $\tau$. This threshold is not universal; its optimal value is domain-specific and depends on both the base model’s calibration and the nature of the task. For instance, highly creative or open-ended tasks might naturally exhibit higher entropy, whereas logical deduction tasks may have a lower baseline entropy. Consequently, $\tau$ must be tuned on a held-out validation set to find the right balance between computational efficiency (favoring a higher $\tau$ to branch less) and exploration (favoring a lower $\tau$ to capture more uncertainties). We detail our tuning process and analyze the sensitivity of our results to $\tau$ in Section 5.
@@
When entropy exceeds the threshold, it indicates a critical decision point where the model is uncertain and multiple reasoning paths may be viable. These are the moments where exploration is most valuable. We employ a rollback-and-branch strategy to precisely locate and expand at uncertainty points. When generating a continuation sequence for beam $i$, we monitor the entropy at each token position. If at any position $t^{\prime}>t$ the entropy $H_{t^{\prime}}^{(i)}$ exceeds $\tau$ during generation, we identify the earliest such position:
$t^{*}=\min\{t^{\prime}\mid t^{\prime}>t\text{ and }H_{t^{\prime}}^{(i)}>\tau\}$ (2)
We then roll back the generation to position $t^{*}$ and create $W$ diverse candidate branches starting precisely from this high-entropy token. Formally, the rollback operation truncates the generated sequence:
$s^{(i)}\leftarrow s^{(i)}_{1:t^{*}-1}$ (3)
and initiates branching at position $t^{*}$, yielding candidate set $\{c_{t^{*},1}^{(i)},\dots,c_{t^{*},W}^{(i)}\}$. This rollback mechanism ensures that branching occurs exactly at the first moment of uncertainty rather than at arbitrarily chosen checkpoints, concentrating search budget precisely where the model’s confidence wavers.
@@ 3.2 Formalizing Selective Expansion
The total number of candidates generated at step $t$ is therefore at most $|\mathcal{P}_{t}|=K+(W-1)|\mathcal{U}_{t}|$. Since $|\mathcal{U}_{t}|\leq K$, this is always less than or equal to the $KW$ candidates produced by standard beam search. The computational complexity of selective expansion scales as $O(|\mathcal{U}_{t}|\cdot W\cdot V)$ for candidate generation, compared to $O(K\cdot W\cdot V)$ for uniform beam search. In practice, since $|\mathcal{U}_{t}|$ is much smaller than $K$ due to the selective nature of branching, EGB achieves substantial computational savings while maintaining search effectiveness.
@@ 3.3 Beam Evaluation
Before proceeding to quality assessment, EGB implements a duplicate detection mechanism to avoid wasting computational resources on scoring identical continuations. Since both uncertain beams and certain beams contribute to the candidate pool $\mathcal{P}_{t}$, identical or near-identical sequences can emerge from different sources, particularly when the model exhibits strong preferences for specific token sequences. Therefore, we employ exact string matching at the token level and remove duplicate branches first, and retain the remaining diverse branches for further evaluation. After collecting candidate continuations, assessing their quality is critical, as it directly influences the overall performance of the base model. A strong feedback model ensures that the generation follows a coherent and logical trajectory while minimizing errors. We employ a Process Reward Model that evaluates the quality of reasoning steps rather than just final outcomes. The PRM assigns scores to partial solutions based on their logical consistency, mathematical accuracy, and adherence to sound reasoning principles. For a candidate continuation $c_{t,j}^{(i)}$, the PRM score is computed as:
@@ 4 Experiments
We evaluate on mathematical reasoning benchmarks of varying difficulty, including AIME Art of Problem Solving Community (2025), MATH-500 Hendrycks et al. (2021), and GSM8K Cobbe et al. (2021). In addition to general math benchmarks, we also include financial math datasets to test our method in a domain-specific setting. We include financial math from CFA mock exams purchased from AnalystPrep AnalystPrep (2024), covering two program levels. CFA exams assess rigorous financial and quantitative reasoning in high-stakes scenarios where errors can cause significant losses. We present more details about the dataset in the Appendix A.
@@
We implement our method on two families of models, Qwen3 Yang et al. (2025) and Llama3 Touvron et al. (2023), at 3 sizes each. For Qwen3 we use Qwen3-1.7B, Qwen3-4B, and Qwen3-8B, all in "non-thinking" mode Yang et al. (2025). For Llama3 we use Llama-3.2-1B-instruct, Llama-3.2-3B -instruct, and Llama-3.1-8B-instruct. Unless otherwise noted, we adopt the recommended decoding settings: for Qwen3, temperature 0.7, top-p 0.8, top-k 20, and min-p 0; for Llama3, temperature 0.6 and top-p 0.9.
@@
Baselines
We benchmark EGB against several strong test-time reasoning methods. Self-consistency Wang et al. (2022) generates multiple independent reasoning paths and aggregates them via majority voting. Self-evaluation-guided stochastic beam search Xie et al. (2023) incorporates model self-evaluation into the beam search process. Step-wise beam search with external PRM feedback Snell et al. (2025) employs deterministic decoding to retain the top-k candidates at each reasoning step.
@@
To balance performance with computational overhead, we set the entropy thresholds $\tau=1.5$ for all Qwen models, and $\tau=0.8$ for the Llama models since they exhibit less uncertainty. For results reported in Table 1, we use a beam size $K$ of 4, a beam width $W$ of 4 for all datasets, and a sample budget of 16 for Self-consistency. Detailed analyses of the threshold values and branch numbers can be found in Section A.2.
@@
All experiments were performed on H100 GPUs with 80GB VRAM. Each experiment was run on one GPU, utilizing 20-70 GB of GPU memory, depending on the model size and sequence length.
@@
Self-consistency 56.89% 44.89% 13.33% 74.23% 84.23%
SEGBS 55.94% 45.45% 7.78% 69.60% 80.67%
Beam Search 58.72% 50.57% 10.00% 73.80% 86.58%
EGB 61.50% 51.70% 11.11% 77.40% 89.46%
@@
Self-consistency 66.67% 56.36% 17.78% 78.60% 92.42%
SEGBS 71.61% 57.39% 17.78% 80.20% 93.18%
Beam Search 72.69% 52.27% 21.11% 81.20% 93.18%
EGB 69.72% 60.80% 18.60% 83.80% 94.01%
@@
Self-consistency 76.56% 59.66% 18.89% 78.40% 92.72%
SEGBS 76.61% 58.52% 17.78% 82.60% 93.10%
Beam Search 76.00% 57.39% 18.89% 82.40% 94.54%
EGB 77.73% 61.36% 22.22% 84.20% 95.45%
@@
Self-consistency 37.56% 31.82% 3.33% 42.80% 52.39%
SEGBS 38.17% 31.82% 11.11% 42.58% 58.21%
Beam Search 43.83% 31.82% 3.33% 45.80% 63.00%
EGB 42.28% 34.09% 3.33% 44.20% 61.92%
@@
Self-consistency 52.17% 40.89% 11.11% 56.39% 72.30%
SEGBS 46.61% 41.48% 11.11% 50.84% 85.24%
Beam Search 51.11% 38.64% 10.00% 55.80% 87.49%
EGB 52.28% 42.61% 11.11% 56.80% 84.69%
@@
Self-consistency 53.83% 50.89% 10.00% 57.84% 90.45%
SEGBS 54.39% 48.18% 11.11% 56.20% 90.65%
Beam Search 56.94% 43.75% 11.11% 60.20% 91.36%
EGB 55.78% 51.14% 11.11% 62.20% 90.67%
Table 1: Performance of EGB compared with beam search, self-evaluation-guided stochastic beam search (SEGBS), self-consistency, and standard decoding on finance and math benchmarks.